# DP-LLM: Runtime Model Adaptation with Dynamic Layer-wise Precision Assignment

### v1

@@ Abstract
How can we effectively handle queries for on-device large language models (LLMs) with varying runtime constraints, such as latency and accuracy? Multi-scale quantization addresses this challenge by enabling memory-efficient runtime model adaptation of LLMs through the overlaying of multiple model variants quantized to different bitwidths. Meanwhile, an important question still remains open-ended: how can models be properly configured to match a target precision or latency? While mixed-precision offers a promising solution, we take this further by leveraging the key observation that the sensitivity of each layer dynamically changes across decoding iterations. Building on this insight, we introduce DP-LLM, a novel mechanism that dynamically assigns precision to each layer based on input values. DP-LLM augments each linear layer in an LLM with a precision selector that determines the bitwidth at runtime using a lightweight error estimator and threshold values learned through fine-tuning. Experimental results across multiple models and benchmarks demonstrate that DP-LLM achieves a superior performance-latency trade-off, outperforming prior approaches.
@@ 1 Introduction
While layer-wise mixed-precision matryoshka ; mq ; hawqv2 can be integrated with multi-scale quantization to address this issue, we identify an overlooked opportunity for further improvement: the sensitivity of each layer dynamically changes across decoding iterations. In other words, a layer that requires a higher bitwidth at certain decoding iterations due to its sensitivity may become less sensitive at other iterations, making a lower bitwidth more suitable. This observation highlights the need for dynamic layer-wise precision assignment at each decoding iteration, rather than static precision assignment.
To leverage this yet unexplored opportunity, we propose Dynamic-Precision LLM (DP-LLM), a model runtime adaptation mechanism that supports dynamic layer-wise precision assignment. DP-LLM (i) assigns high and low precisions and a threshold of relative error for each layer to determine which precision to use and (ii) employs a lightweight precision selector to each layer that estimates the relative error with minimal latency overhead at each decoding iteration. Through extensive experiments on various datasets and downstream tasks, we demonstrate that DP-LLM achieves a superior performance-latency trade-off compared to prior methods.
@@
  • We find that layer-wise sensitivity to quantization is not a static property, but rather changes dynamically with each decoding iteration (token by token).
  • We devise a mechanism that sets a threshold for each layer to dynamically determine which precision to assign.
  • We devise a lightweight precision selector that can effectively estimate the relative error to compare it against the assigned threshold.
@@ 2.2 Runtime Model Adaptation for LLMs
Different queries in DNN serving often have varying runtime constraints, such as accuracy, latency, or energy requirements. Edge devices are no exception, as the volatility in available resources and the variety of downstream tasks dynamically affect the runtime constraints. An effective approach to address these varying constraints is to maintain a set of models with different trade-offs between accuracy and overhead (e.g., latency, throughput, and energy), which we refer to as adaptation set, and select the most suitable one based on the runtime requirements for each query model-adapt-3 ; adabits ; model-adapt-1 ; model-adapt-2 ; bitmixer .
Figure 1 illustrates an example of runtime model adaptation. The adaptation set is configured under a given memory budget, and consists of models with different precisions. The gap between the varying query-wise resource budget under QoS (quality-of-service) constraints, or QoS budget for brevity, and fluctuating system utilization creates volatility in the LLM inference slack, and at any given time, the model with the precision that best fills the slack is selected. For example, when the QoS budget is relaxed and system utilization is low, a higher-precision model such as a 7-bit variant is preferred. Conversely, when the QoS budget is tight and/or system utilization is high—leaving limited slack—a lower-precision model (e.g., 3-bit or 4-bit) may be selected.
@@ 2.4 Opportunities for Dynamic Layer-wise Mixed-Precision
One aspect that has not received much attention regarding layer-wise mixed-precision in the context of LLM inference is the dynamically changing sensitivity of each layer during the decoding phase. That is, layers that are sensitive at certain decoding iterations may become less sensitive at others, and vice versa. However, prior works assign the bitwidth of layers statically: once a specific bitwidth is assigned to a layer, it remains fixed throughout the decoding phase, failing to account for this dynamic behavior matryoshka ; mq ; hawqv2 ; aqdnn ; hawq .
Figure 3(a) demonstrates such dynamically changing sensitivity of layers. This figure shows the distribution of the top 20% most sensitive layers at each decoding iteration when decoding with a Llama-3-8B model, using the first sample from the C4 dataset with only the first half layers depicted. Assuming a scenario where 3-bit and 4-bit precision are used to configure a mixed-precision model, the sensitivity of each layer is defined as the decrease in perplexity when applying 4-bit precision to that layer while keeping the rest at 3-bit. The distribution is highly irregular, which calls for dynamic bitwidth assignment for each layer at each decoding iteration. Figure 2(c) depicts our proposed dynamic layer-wise mixed-precision.
Figure 3: (a) Sensitivity of different layers at each decoding iteration. (b) Perplexity trend of different precision assignment schemes.
Figure 3(b) shows the perplexity of a dynamic layer-wise mixed-precision model that dynamically chooses between 3-bit and 4-bit for each layer based on sensitivity at each decoding iteration, using the same sensitivity analysis method as in Figure 3(a). While the exact implementation of this scheme is practically infeasible—since figuring out sensitivity at runtime is not possible—it serves as an indicator of the potential improvements achievable through the concept of dynamic layer-wise mixed-precision. This scheme is compared to static layer-wise mixed-precision, which assigns bitwidth to layers statically based on their average sensitivity. Dynamic layer-wise mixed-precision demonstrates substantial improvements over the static approach, highlighting the importance of leveraging the dynamic behavior of layer-wise sensitivity.
@@ 3 DP-LLM Overview
Leveraging the dynamic nature of layer-wise sensitivity discussed in Section 2.4, we introduce Dynamic-Precision LLM (DP-LLM), a novel approach that employs a dynamic, layer-wise mixed-precision scheme for runtime model adaptation in LLMs. Figure 4 presents an overview of DP-LLM. Instead of preconfiguring the adaptation set offline, DP-LLM dynamically configures the model according to the target precision at runtime. Specifically, each layer dynamically selects either $h$-bit weights ($W_{h}$) or $l$-bit weights ($W_{l}$), where $h>l$, based on input values at each decoding iteration. To achieve this, DP-LLM augments each linear operation (essentially a GEMV operation when decoding with a batch size of 1) in the Transformer block with a precision selector. This precision selector processes the input vector and, based on its values, determines the appropriate precision.
@@
The main insight behind designing the precision selector is that the norm of the difference in the GEMV output between using $W_{h}$ and $W_{l}$, or $\Delta Wx||$ (where $\Delta W=W_{h}-W_{l}$ and $x$ is the input vector), can serve as an effective indicator determining which precision to use. We refer to this difference in magnitude as relative error. Intuitively, it is preferable to use $W_{l}$ for inputs that result in small relative errors, while $W_{h}$ should be prioritized for inputs likely to produce larger relative errors.
Such an approach necessitates addressing two key research questions. The first challenge is determining a threshold $T$, along with $h$ and $l$ for each layer: inputs whose relative error exceeds $T$ use the higher-precision weights $W_{h}$, while those below $T$ use the lower-precision weights $W_{l}$. These $T$ values must be carefully determined, considering the impact of each layer on the end-to-end loss. Additionally, the $T$ values should be configured in a way that ensures the average bitwidth across the model closely matches the target precision. Second, we need a scheme to approximate the relative error with minimal computation overheads. Since relative error guides our bitwidth selection, we must be able to estimate it efficiently. The exact computation of relative error is infeasible at runtime, as $\Delta W$ has the same precision and dimension as of the dequantized weights.
Section 4 explains how DP-LLM addresses the first research question, while Section 5 explains how DP-LLM tackles the second research question.
4 Layer-wise Threshold Assignment
Figure 5: Overview of layer-wise threshold assignment
The process of determining $T$ values for each layer is divided into three phases: (1) Layer-wise maximum precision selection (2) Layer-wise average precision assignment and (3) Average precision-to-threshold translation. Figure 5 provides an overview of these phases.
@@ B $\leftarrow{}$ Find maximum precision for each layer $i$ using S, $M_{max}$
Fine-tune $\{p_{i}\}$ for $\mathcal{M}$ with $D$
@@
  ${\Delta}W_{i}$ $\leftarrow{}$ $W_{i,h}-W_{i,l}$
@@
Algorithm 1 Layer-wise Threshold Assignment of DP-LLM
@@
DP-LLM adapts to the given memory budget by first selecting the max precisions for each layer ($B[i]$ in Algorithm 1) before determining the average precisions. Inspired by mq ; hawqv2 ; sqllm , we use the second-order Taylor expansion on the loss function as the static sensitivity metric and form an integer programming problem to select maximum precisions. The formulation is explained in Appendix A. The range of $p_{i}$ is adjusted to match the selected max precision of each layer during fine-tuning. Figure 5(a) corresponds to this phase.
@@
For dynamic precision selection, DP-LLM determines average precision, $p$, for each layer. The average precision represents the expectation of selected precisions throughout the decoding phase for each layer. To construct an average precision, we select $l={\lfloor}p{\rfloor}$ and $h={\lceil}p{\rceil}$ precision weights as our bases. For example, a layer with a $p$ value of 3.2 has approximately 80% probability of selecting 3-bit and 20% of selecting 4-bit weight at each decoding iteration. An ablation study for selecting different precisions as bases is in Appendix C.3.
DP-LLM assigns a distinct $p$ value to each layer considering the sensitivity of each layer. For example in Figure 5(b), the target precision of 3.9 does not necessarily imply that $p=3.9$ for all layers. Instead, different layers can have varying $p$ values (and varying $l$ and $h$ accordingly), as long as the overall average results in the target precision.
DP-LLM parameterizes $p$ values for each layer and performs fine-tuning to optimize these values with respect to the end-to-end loss. Specifically, during the fine-tuning process, the original linear layer $y=Wx$ is substituted with $y=rW_{l}x+(1-r)W_{h}x$}, where $l={\lfloor}p{\rfloor}$, $h={\lceil}p{\rceil}$, and r=1(pl)r=1-(p-l). This formulation, however, naturally causes the $p$ values to collapse to $h$ (and eventually to the highest precision possible), as using high bitwidth weights is generally more advantageous in terms of minimizing the loss. This outcome is undesirable, as it deviates from the target precision. To address this, we introduce a regularization term into the loss function to ensure that the overall average of $p$ values across all layers aligns with the target precision. Equation 1 shows the modified loss function, where $M_{i}$ is the number of parameters in layer $i$, $b_{targ}$ is the target precision, and $\alpha$ is a hyperparameter controlling the regularization strength.
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}^{\prime}=\mathcal{L}+\alpha{\left({\left.{\displaystyle\sum_{i}^{N}{p_{i}M_{i}}}\middle/{\displaystyle\sum_{i}^{N}{M_{i}}}\right.}~-~b_{targ}\right)^{2}}$} (1)
This fine-tuning process is conducted on a small calibration dataset with only a few iterations. In addition, the $p$ values are the only parameters updated during the fine-tuning process. Thus, our approach keeps both memory and computational costs of fine-tuning low. The training costs are reported in Appendix B.3.
@@
Once the average precisions have been assigned to each layer, they are translated into threshold values through statistical analysis using the calibration dataset—the same data samples used for average precision assignment. Specifically, DP-LLM leverages the distribution of relative errors from the calibration dataset to approximate the distribution of runtime inputs. For each layer, ${\Delta}W_{i}x||$ is calculated during the calibration stage. Then, among the distribution of relative errors, the $r_{i}$-quantile value is selected as the threshold, where $r_{i}=1-(p_{i}-l)$.
@@ 5.1 Hybrid Approach for Relative Error Estimation
DP-LLM adopts a hybrid approach for relative error estimation, selecting one of two methods based on whether a strong linear relationship exists between the input vector’s norm $x||$ and the relative error $\Delta Wx||$ for each layer: Linear regression-based or random projection-based estimation.
For layers with strong correlation—specifically, those where the coefficient of determination ($R^{2}$) exceeds 0.9—DP-LLM employs simple linear regression to predict the relative error. In this case, $\Delta Wx||$ is estimated using a simple linear function with $x||$, given by $x||\times\alpha+\beta$. This approach incurs near-zero latency and GPU memory overhead. Notably, approximately half of the layers satisfy this property.
For layers without a strong correlation, DP-LLM utilizes random projection leveraging Johnson-Lindenstrauss Lemma jl (JL Lemma) to efficiently perform relative error estimation in a low-dimensional space. JL Lemma states that the probability of satisfying Inequality 2 for any $n$-dimensional vector $x$ is $1-\delta$ when $A$ is a $k\times n$ matrix sampled from a normal distribution $A_{ij}\sim{1\over{\sqrt{k}}}N(0,1)$ and $k=\mathcal{O}(\epsilon^{-2}\log(\delta^{-1}))$. In other words, estimating a vector’s norm with a degree of confidence can be done by calculating the norm of the multiplication of a randomized matrix and the vector. In our case of estimating the input token’s error, we need to estimate the norm of $\Delta Wx$. Therefore, we plug in $\Delta Wx$ to the lemma to get Inequality 3.
@@
By precomputing $G=A\Delta W$, the relative error can be estimated at runtime by performing a GEMV operation on $G$ and $x$, which is an $\mathcal{O}(nk)$ operation. In our further experiment setups, we use $k=64$ for every linear projection, which limits the relative error estimation difference within 15% with 91% confidence when measured empirically using the C4 dataset. The error difference can be further reduced by calibrating $G$ with the distribution of $x$. Using the calibration set, we use gradient descent on $G$ to match the input distribution. This procedure can be performed in parallel for each layer, and only takes a small fraction of time (<10 seconds per layer).

### v3

@@ Abstract
How can we effectively handle queries for on-device large language models (LLMs) with varying runtime constraints, such as latency and accuracy? Multi-scale quantization addresses this challenge by enabling memory-efficient runtime model adaptation of LLMs through the overlaying of multiple model variants quantized to different bitwidths. Meanwhile, an important question still remains open-ended: how can models be properly configured to match a target precision or latency? While mixed-precision offers a promising solution, we take this further by leveraging the key observation that the sensitivity of each layer dynamically changes across decoding steps. Building on this insight, we introduce DP-LLM, a novel mechanism that dynamically assigns precision to each layer based on input values. Experimental results across multiple models and benchmarks demonstrate that DP-LLM achieves a superior performance-latency trade-off, outperforming prior approaches.
@@ 1 Introduction
While layer-wise mixed-precision matryoshka ; mq ; hawqv2 can be integrated with multi-scale quantization to address this issue, we identify an overlooked opportunity for further improvement: the sensitivity of each layer dynamically changes across decoding steps. In other words, a layer that requires a higher bitwidth at certain decoding steps due to its sensitivity may become less sensitive at other steps, making a lower bitwidth more suitable. This observation highlights the need for dynamic layer-wise precision assignment at each decoding step, rather than static precision assignment.
To leverage this yet unexplored opportunity, we propose Dynamic-Precision LLM (DP-LLM), a model runtime adaptation mechanism that supports dynamic layer-wise precision assignment. DP-LLM defines a proxy metric of quantization sensitivity, called relative error, and determines for each layer, at offline, when to use high- or low-bit precision based on the magnitude of this proxy metric. At runtime, DP-LLM efficiently estimates the relative error and selects the appropriate precision for each layer. Through extensive experiments on various datasets and downstream tasks, we demonstrate that DP-LLM achieves a superior performance-latency trade-off compared to prior methods.
@@
  • We find that layer-wise sensitivity to quantization is not a static property, but rather changes dynamically with each decoding step (token by token).
  • We define a proxy metric, termed relative error, which serves as a criterion for precision assignment, and design a mechanism that learns, for each layer, how to determine which precision to apply based on the value of this metric.
  • We develop a lightweight precision selector that efficiently estimates the relative error and selects the appropriate precision at runtime.
@@ 2.2 Runtime Model Adaptation for LLMs
Different queries in DNN serving often have varying runtime constraints, such as accuracy, latency, or energy requirements. Edge devices are no exception, as the volatility in available resources memutil ; approxdet and the variety of downstream tasks llminfsys ; distserve dynamically affect the runtime constraints. An effective approach to address these varying constraints is to maintain a set of models with different trade-offs between accuracy and overhead (e.g., latency, throughput, and energy), which we refer to as adaptation set, and select the most suitable one based on the runtime requirements for each query model-adapt-3 ; adabits ; model-adapt-1 ; model-adapt-2 ; bitmixer .
Figure 1 illustrates an example of runtime model adaptation. The adaptation set is configured under a given memory budget, and consists of models with different precisions. The gap between the varying query-wise resource budget under QoS (quality-of-service) constraints, or QoS budget for brevity, and the fluctuating system utilization creates volatility in the LLM inference slack, and at any given time, the model with the precision that best fills the slack is selected. For example, when the QoS budget is relaxed and the system utilization is low, a higher-precision model such as a 7-bit variant is preferred. Conversely, when the QoS budget is tight and/or the system utilization is high—leaving limited slack—a lower-precision model (e.g., 3-bit or 4-bit) may be selected.
@@ 2.4 Opportunities for Dynamic Layer-wise Mixed-Precision
One aspect that has not received much attention regarding layer-wise mixed-precision in the context of LLM inference is the dynamically changing sensitivity of each layer during the decoding phase. That is, layers that are sensitive at certain decoding steps may become less sensitive at others, and vice versa. However, prior works assign the bitwidth of layers statically: once a specific bitwidth is assigned to a layer, it remains fixed throughout the decoding phase, failing to account for this dynamic behavior matryoshka ; mq ; hawqv2 ; aqdnn ; hawq .
Figure 3(a) demonstrates such dynamically changing sensitivity of layers. This figure shows the distribution of the top 20% most sensitive layers at each decoding step when decoding with a Llama-3-8B model, using the first sample from the C4 dataset with only the first half layers depicted. Assuming a scenario where 3-bit and 4-bit precision are used to configure a mixed-precision model, the sensitivity of each layer is defined as the decrease in perplexity when applying 4-bit precision to that layer while keeping the rest at 3-bit. The distribution is highly irregular, which calls for dynamic bitwidth assignment for each layer at each decoding step. Figure 2(c) depicts our proposed dynamic layer-wise mixed-precision.
Figure 3: (a) Sensitivity of different layers at each decoding step. (b) Perplexity trend of different precision assignment schemes.
Figure 3(b) shows the perplexity of a dynamic layer-wise mixed-precision model that dynamically chooses between 3-bit and 4-bit for each layer based on sensitivity at each decoding step, using the same sensitivity analysis method as in Figure 3(a). While the exact implementation of this scheme is practically infeasible—since figuring out sensitivity at runtime is not possible—it serves as an indicator of the potential improvements achievable through the concept of dynamic layer-wise mixed-precision. This scheme is compared to a static layer-wise mixed-precision scheme, which assigns bitwidth to layers statically based on their average sensitivity. Dynamic layer-wise mixed-precision demonstrates substantial improvements over the static approach, highlighting the importance of leveraging the dynamic behavior of layer-wise sensitivity.
@@ 3 DP-LLM Overview
Leveraging the dynamic nature of layer-wise sensitivity discussed in Section 2.4, we introduce Dynamic-Precision LLM (DP-LLM), a novel approach that employs a dynamic, layer-wise mixed-precision scheme for runtime model adaptation in LLMs. Figure 4 presents an overview of DP-LLM. Instead of preconfiguring the adaptation set offline, DP-LLM dynamically configures the model according to the target precision at runtime. Specifically, at offline, DP-LLM assigns each layer a candidate precision set, a set of precision levels that can be selected at runtime based on input values in each decoding step. To limit complexity, each candidate set consists of two precision levels, $h$-bit and $l$-bit, where $h>l$. To enable such runtime selection, DP-LLM augments each linear operation (essentially a GEMV operation when decoding with a batch size of 1) in the Transformer block with a precision selector. This precision selector processes the input vector and, based on its values, determines the appropriate precision.
@@
The main insight behind designing the precision selector is that the norm of the difference in the GEMV output between using $W_{h}$ ($h$-bit weight) and $W_{l}$ ($l$-bit weight), or $\Delta Wx||$ (where $\Delta W=W_{h}-W_{l}$ and $x$ is the input vector), can serve as an effective indicator determining which precision to use. We refer to this difference in magnitude as relative error. Intuitively, it is preferable to use $W_{l}$ for inputs that result in small relative errors, while $W_{h}$ should be prioritized for inputs that are likely to produce larger relative errors.
Such an approach necessitates addressing two key challenges. The first challenge is determining a threshold $T$, along with $h$ and $l$ for each layer: inputs whose relative error exceeds $T$ use the higher-precision weights $W_{h}$, while those below $T$ use the lower-precision weights $W_{l}$. These $T$ values must be carefully determined, considering the impact of each layer on the end-to-end loss. Additionally, the $T$ values should be configured in a way that ensures the average bitwidth across the model closely matches the target precision. The second challenge lies in devising an efficient scheme to approximate the relative error with minimal computational overhead. While relative error guides our bitwidth selection, its exact computation is infeasible at runtime, as it would require performing an additional GEMV operation.
Section 4 explains how DP-LLM addresses the first challenge, while Section 5 explains how DP-LLM tackles the second.
4 Layer-wise Candidate Precision Set and Threshold Assignment
Figure 5: Overview of layer-wise candidate precision set and threshold assignment
The process of determining candidate precision set and threshold (i.e., $h$, $l$ and $T$) for each layer is divided into three phases: (1) Layer-wise maximum precision selection (2) Layer-wise average precision assignment and (3) Average precision-to-threshold translation. Figure 5 provides an overview of these phases, while Algorithm 1 provides the overall procedures.
@@ 4 Layer-wise Candidate Precision Set and Threshold Assignment
Fine-tune $\{p_{i}\}$ for $\mathcal{M}$ with $D$ using Equation 1 as loss function
@@
  ${\Delta}W_{i}$ $\leftarrow{}$ $W_{i,h}-W_{i,l}$ where $l=\lfloor p_{i}\rfloor$ and $h=\lceil p_{i}\rceil$
  Initialize err_list as empty list
@@
Algorithm 1 Layer-wise Candidate Precision Set and Threshold Assignment of DP-LLM
@@
DP-LLM adapts to the given memory budget by first selecting the maximum precision for each layer (List $B$ in Algorithm 1) before determining the average precisions. For example, in Figure 5(a), the maximum precisions for Layers 1, 2, and 3 are set to 5, 4, and 5, respectively. Inspired by mq ; hawqv2 ; sqllm , we use the second-order Taylor expansion on the loss function as the static sensitivity metric and form an integer programming problem to select maximum precisions. The detailed formulation is provided in Appendix A.
@@
For dynamic precision selection, DP-LLM determines the average precision, $p$, for each layer. The average precision represents the expectation of selected precisions throughout the decoding phase for each layer. Once $p$ is determined, the corresponding candidate precision set is defined as $l={\lfloor}p{\rfloor}$ and $h={\lceil}p{\rceil}$. For example, a layer with a $p$ value of 3.2 would have a candidate precision set of 3-bit and 4-bit, and is expected to select 3-bit weights for 80% of the decoding steps and 4-bit weights for 20% of them throughout the decoding phase.
DP-LLM assigns a distinct $p$ value to each layer considering the sensitivity of each layer. For example, in Figure 5(b), the target precision of 3.9 does not necessarily imply that $p=3.9$ for all layers. Instead, different layers can have varying $p$ values (and varying $l$ and $h$ accordingly), as long as the overall average results in the target precision.
DP-LLM parameterizes $p$ values and performs fine-tuning to optimize these values with respect to the end-to-end loss. Specifically, during the fine-tuning process, the original linear layer $y=Wx$ is substituted with $y=rW_{l}x+(1-r)W_{h}x$}, where $l={\lfloor}p{\rfloor}$, $h={\lceil}p{\rceil}$, and r=1(pl)r=1-(p-l). This formulation, however, naturally causes the $p$ values to collapse to $h$ (and eventually to the highest precision possible), as using high bitwidth weights is generally more advantageous in terms of minimizing the loss. This outcome is undesirable, as it deviates from the target precision. To address this, we introduce a regularization term into the loss function to ensure that the overall average of $p$ values across all layers aligns with the target precision. Equation 1 shows the modified loss function, where $M_{i}$ is the number of parameters in layer $i$, $b_{targ}$ is the target precision, and $\alpha$ is a hyperparameter controlling the regularization strength.
$\mathcal{L}^{\prime}=\mathcal{L}+\alpha{\left({\left.{\displaystyle\sum_{i}^{N}{p_{i}M_{i}}}\middle/{\displaystyle\sum_{i}^{N}{M_{i}}}\right.}~-~b_{targ}\right)^{2}}$ (1)
This fine-tuning process is conducted on a small calibration dataset with only a few iterations. In addition, the $p$ values are the only parameters updated during the fine-tuning process. Thus, both memory and computational costs are kept at a manageable level. An analysis of the fine-tuning cost is provided in Appendix B.3.
@@
Once the average precisions have been assigned to each layer, they are translated into threshold values through statistical analysis using the calibration dataset—the same data samples used for average precision assignment. Specifically, DP-LLM leverages the distribution of relative errors from the calibration dataset to approximate the distribution of runtime inputs. For each layer, ${\Delta}W_{i}x||$ is calculated during the calibration stage. Then, among the distribution of relative errors, the $r_{i}$-quantile value is selected as the threshold, where $r_{i}=1-(p_{i}-l)$. Figure 5(c) visualizes this process.
@@ 5.1 Hybrid Approach for Relative Error Estimation
DP-LLM adopts a hybrid approach for relative error estimation, where each layer selects one of two methods: linear regression-based or random projection-based estimation. The former is highly lightweight but relies on a strong linear relationship between the input vector norm $x||$ and the relative error $\Delta Wx||$ for accurate estimation. Hence, it is applied only to layers exhibiting such a relationship. For the remaining layers, the latter method is used; although less lightweight, it performs robustly across diverse scenarios. To determine the strength of the linear relationship, the coefficient of determination ($R^{2}$) between the input vector norm and the relative error for each layer is computed using the calibration set, and compared against a hyperparameter $R^{2}_{th}$(which is set to 0.9 in our further experiment setups) at offline.
Linear Regression-Based Estimation
[…]
Random Projection-Based Estimation
For layers without a strong correlation, DP-LLM utilizes random projection leveraging the Johnson-Lindenstrauss Lemma jl (JL Lemma) to efficiently perform relative error estimation in a low-dimensional space. JL Lemma states that the probability of satisfying Inequality 2 for any $n$-dimensional vector $x$ is $1-\delta$ when $A$ is a $k\times n$ matrix sampled from a normal distribution $A_{ij}\sim{1\over{\sqrt{k}}}N(0,1)$ and $k=\mathcal{O}(\epsilon^{-2}\log(\delta^{-1}))$. In other words, estimating a vector’s norm with a degree of confidence can be done by calculating the norm of the multiplication of a randomized matrix and the vector. In our case of estimating the input token’s error, we need to estimate the norm of $\Delta Wx$. Therefore, we plug in $\Delta Wx$ to the lemma to get Inequality 3.
@@
By precomputing $G=A\Delta W$, the relative error can be estimated at runtime by performing a small GEMV operation on $G$ and $x$, which is an $\mathcal{O}(nk)$ operation. In our further experiment setups, we use $k=64$ for every linear projection, which limits the relative error estimation difference within 15% with 91% confidence when measured empirically using the C4 dataset. The error difference can be further reduced by calibrating $G$ with the distribution of $x$. Using the calibration set, we tune $G$ to match the input distribution at offline. This procedure can be performed in parallel for each layer, and only takes a small fraction of time (<10 seconds per layer).